# The multipartite entanglement classes of a multiport beam-splitter

### v3

@@ III.1 Superpositions of number states as input
where the submatrices $M_{\overline{k}}$ are indexed according to their first column $k$ with decimal expression $n_{2}n_{3}\ldots n_{m}$; notice that $k$ is then a multiple of $N+1$. This notation is to take into account the repetitions of submatrices that occur for $m>3$, as well as identifying null submatrices and allowing possible future generalizations for the infinite superposition case. Moreover, let $||\vec{k}||=n_{2}+n_{3}+\ldots+n_{m}$. The submatrices $M_{\overline{k}}$ are Hankel matrices horn of a special form,
@@
$||\vec{k}||=2,$
and an arbitrary submatrix with $||\vec{k}||=n$ is given by
@@
Notice that if $||\vec{k}||>N$, the submatrix $M_{\overline{k}}$ is a null matrix. Let $A=A^{(N-1)}A^{(N-2)}\ldots A^{(1)}A^{(0)}$, where $A^{(k)}=\mathcal{I}+\sum_{n=k}^{N-1}\lambda_{n}|n\rangle\langle N-k|$, with $\mathcal{I}$ denoting the identity operator and $\lambda_{n}=-h_{n}/h_{N}$. The matrices representing these operators in number basis are given by
@@
while the various $M_{\overline{k}}$ with $||\vec{k}||=1$ are mapped into
@@
maps the various $A^{(1)}M_{\overline{k}}$ with $||\vec{k}||=1$ into
@@
and the various $M_{\overline{k}}$ with $||\vec{k}||=2$ into
@@
while the other submatrices are unchanged. It is easy to see then that the overall effect of $A_{1}|\Phi\rangle$ on the various submatrices such that $||\vec{k}||=n$ is
@@ III.2 Superpositions of coherent states as input
and we have shown that $|\Psi\rangle$ is SLOCC-equivalent to the $m$-partite $GHZ$ state $|GHZ_{(r)}\rangle$. $\Box$
@@ IV Comparison between the three scenarios
By adjoint reduced density matrices it is understood the density matrices obtained through discarding (partial trace) of one of the local subsystems $\mathcal{H}_{k}$ of the global system $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\ldots\otimes\mathcal{H}_{m}$. Denoting by $a_{k}$ the number of product states in the range of the adjoint reduced density matrix related to $\mathcal{H}_{k}$, it is a consequence of Theorem 1 above that if two states are SLOCC-equivalent, then the array of values $[a_{1},a_{2},\ldots,a_{m}]$ of these states must be equal. Or, equivalently, if any value $a_{k}$ differs for these states, then they are SLOCC-inequivalent.
@@
It is straightforward then that each representative is SLOCC-inequivalent to the other, even if they have the same Schmidt rank. For example, the Schmidt rank of $|GHZ_{(M)}\rangle$, $|\Psi_{M-1}\rangle$ and $|\Phi_{M-2}\rangle+|M-1\rangle^{\otimes m}$ is $M$, but each is respectively in $[M,M,\ldots,M]$, $[1,1,\ldots,1]$ and $[2,2,\ldots,2]$, being thus inequivalent by SLOCC.

### v4

@@ III.1 Superpositions of number states as input
where the submatrices $M_{\overline{k}}$ are indexed according to their first column $k$ with decimal expression $n_{2}n_{3}\ldots n_{m}$; notice that $k$ is then a multiple of $N+1$. This notation is to take into account the repetitions of submatrices that occur for $m>3$, as well as identifying null submatrices and allowing possible future generalizations for the infinite superposition case. Moreover, let $\vec{k}||=n_{2}+n_{3}+\ldots+n_{m}$. The submatrices $M_{\overline{k}}$ are Hankel matrices horn of a special form,
@@
[…]
and an arbitrary submatrix with $\vec{k}||=n$ is given by
@@
Notice that if $\vec{k}||>N$, the submatrix $M_{\overline{k}}$ is a null matrix. Let $A=A^{(N-1)}A^{(N-2)}\ldots A^{(1)}A^{(0)}$, where $A^{(k)}=\mathcal{I}+\sum_{n=k}^{N-1}\lambda_{n}|n\rangle\langle N-k|$, with $\mathcal{I}$ denoting the identity operator and $\lambda_{n}=-h_{n}/h_{N}$. The matrices representing these operators in number basis are given by
@@
while the various $M_{\overline{k}}$ with $\vec{k}||=1$ are mapped into
@@
maps the various $A^{(1)}M_{\overline{k}}$ with $\vec{k}||=1$ into
@@
and the various $M_{\overline{k}}$ with $\vec{k}||=2$ into
@@
while the other submatrices are unchanged. It is easy to see then that the overall effect of $A_{1}|\Phi\rangle$ on the various submatrices such that $\vec{k}||=n$ is
@@ III.2 Superpositions of coherent states as input
and we have shown that $|\Psi\rangle$ is SLOCC-equivalent to the $m$-partite GHZGHZ state $|GHZ_{(r)}\rangle$. $\Box$
@@ IV Comparison between the three scenarios
By adjoint reduced density matrices it is understood the density matrices obtained through discarding (partial trace) of one of the local subsystems $\mathcal{H}_{k}$ of the global system $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\ldots\otimes\mathcal{H}_{m}$. Denoting by $a_{k}$ the number of product states in the range of the adjoint reduced density matrix related to $\mathcal{H}_{k}$, it is a consequence of Theorem 1 above that if two states are SLOCC-equivalent, then the array of values $a_{1},a_{2},\ldots,a_{m}]$ of these states must be equal. Or, equivalently, if any value $a_{k}$ differs for these states, then they are SLOCC-inequivalent.
@@
It is straightforward then that each representative is SLOCC-inequivalent to the other, even if they have the same Schmidt rank. For example, the Schmidt rank of $|GHZ_{(M)}\rangle$, $|\Psi_{M-1}\rangle$ and $|\Phi_{M-2}\rangle+|M-1\rangle^{\otimes m}$ is $M$, but each is respectively in $M,M,\ldots,M]$, $1,1,\ldots,1]$ and $2,2,\ldots,2]$, being thus inequivalent by SLOCC.